# Optimal Spectral Transitions in High-Dimensional Multi-Index Models

### v1

@@ 1 Introduction
A popular model to study learning problems in statistical, computer science and machine learning is the multi-index model, where the target function depends on a low-dimensional subspace of the covariates. In this problem, one aim at identifying the $p$-dimensional linear subspace spanned by a family of orthonormal vectors ${\bm{w}}_{\star,1},\dots,{\bm{w}}_{\star,p}\in\mathbb{R}^{d}$ from $n$ independent observations $({\bm{x}}_{i},y_{i})$ from the model:
@@
A classical problem in statistics Friedman and Stuetzle (1981); Yuan (2011); Babichev and Bach (2018), the multi-index model has recently gained in popularity in the machine learning community as a generative model for supervised learning data where the labels only depend on an underlying low-dimensional latent subspace of the covariates Aubin et al. (2018); Damian et al. (2022); Dandi et al. (2024a).
Of particular interest to this work are spectral methods, which play a fundamental role in machine learning by offering an efficient and computationally tractable approach to extracting meaningful structure from high-dimensional noisy data. A paradigmatic example is the Baik–Ben Arous–Péché (BBP) transition Baik et al. (2005), where the leading eigenvalue of a matrix correlates with the hidden signal — a phenomenon that is ubiquitous in machine learning theory. Beyond their practical utility, spectral methods often serve as a starting point for more advanced approaches, including iterative and nonlinear techniques. This leads us to the central question of this paper:
@@
While the optimal spectral method for single-index models are well understood Luo et al. (2019); Mondelli and Montanari (2018); Maillard et al. (2022); Mondelli et al. (2022), and their optimality in terms of weak recovery threshold established Barbier et al. (2019); Mondelli and Montanari (2018) their counterparts for multi-index models remain largely unexplored. This gap is particularly important, as multi-index models serve as a natural testbed for studying the computational complexity of feature learning in modern machine learning, and have attracted much attention recently Abbe et al. (2023); Damian et al. (2022); Dandi et al. (2024a); Arnaboldi et al. (2023); Collins-Woodfin et al. (2024); Berthier et al. (2024); Simsek et al. (2024).
@@
In this work, we step up to this challenge by constructing optimal spectral methods for multi-index models. We introduce and analyze spectral algorithms based on a linearized message-passing framework, specifically tailored to this setting. Our main contribution is to present two such constructions, establish the reconstruction threshold for these methods and to show that they achieve the provably optimal threshold for weak recovery among efficient algorithms Troiani et al. (2024).
@@
  • Recently, multi-index models have become a proxy model for studying non-convex optimization Veiga et al. (2022); Arnaboldi et al. (2023); Collins-Woodfin et al. (2024). Arous et al. (2021) has shown that the sample complexity of one-pass SGD for single-index model is governed by the information exponent of the link function. This analysis was generalized in several directions, such as to other architectures Berthier et al. (2024), larger batch sizes Arnaboldi et al. (2024a) and to overparametrised models Arnaboldi et al. (2024c). A similar notion, known as the leap exponent, was introduced for multi-index models, where it was shown that different index subspaces might be learned hierarchically according to their interaction Abbe et al. (2022, 2023); Bietti et al. (2023); Mousavi-Hosseini et al. (2024). The picture was found to be different for batch algorithms exploiting correlations in the data Dandi et al. (2024b); Arnaboldi et al. (2024b); Lee et al. (2024), achieving a sample complexity closer to optimal first order methods Barbier et al. (2019); Damian et al. (2024); Troiani et al. (2024). Troiani et al. (2024) in particular, provided optimal asymptotic thresholds for weak recovery within the class of first order methods.
@@
  • Spectral methods are widely employed as a warm start strategy for initializing other algorithms, in particular for iterative schemes (such as gradient descent) for which random initialization is a fixed point. Relevant to this work is the class of approximate message passing (AMP) algorithms, which have garnered significant interest in the high-dimensional statistics and machine learning communities over the past decade Donoho et al. (2009); Bayati and Montanari (2011); Rangan (2011); Fletcher and Rangan (2018). AMP for multi-index models was discussed in Aubin et al. (2018); Troiani et al. (2024). Spectral initialization for AMP in the context of single-index models has been studied by Mondelli and Venkataramanan (2021).
@@
  • The interplay between AMP and spectral methods has been extensively studied in the literature, see for example Saade et al. (2014); Lesieur et al. (2017); Aubin et al. (2019); Mondelli and Montanari (2018); Mondelli et al. (2022); Maillard et al. (2022); Venkataramanan et al. (2022). In particular the idea of using message passing algorithm to derive spectral method has been very successful, leading to the non-backtracking matrix Krzakala et al. (2013) and more recently to the Kikuchi hierarchy Wein et al. (2019); Hsieh et al. (2023), and to non-linear optimal spectral methods for matrix and tensor factorization Lesieur et al. (2017); Perry et al. (2018); Guionnet et al. (2023); Pak et al. (2024).
@@ 2 Setting and Notations
To enhance readability, we adopt the following consistent notations through the paper: scalar quantities are denoted using non-bold font (e.g., $a$ or $A$), vectors are represented in bold lowercase letters (e.g., $\bm{a}$), matrices are written in bold uppercase letters (e.g., $\bm{A}$), and tensors are depicted using underlined bold uppercase letters (e.g., ${\underline{\bm{A}}}$). We further differentiate random variables depending on the noise with non-italic font (e.g, ${\mathrm{a}},\bm{\mathbf{a}},\bm{\mathbf{A}},{\underline{\bm{\mathbf{A}}}}$). We denote by $\langle\bm{a},\bm{b}\rangle$ the standard Euclidean scalar product, $\|\bm{a}\|$ the $\ell^{2}$-norm of a vector $\bm{a}$, $\|\bm{A}\|_{\mathrm{op}}$ the operator norm of a matrix $\bm{A}$ and $\|\bm{A}\|_{\mathrm{F}}$ its Frobenius norm. Given $n\in\mathbb{N}$, we use the shorthand $\llbracket n\rrbracket=\{1,\dots,n\}$. We denote $\mathbb{S}_{+}^{p}$ the cone of positive semi-definite $p\times p$ matrices. $(x)_{+}:=\max(0,x)$.
@@
and the labels’ marginal distribution
$\mathsf{Z}(y)=\mathbb{E}_{{\bm{W}}_{\star}}\mathbb{E}_{\bm{\mathbf{x}}\sim\mathsf{N}({\bm{0}},d^{-1}{\bm{I}}_{d})}[\mathsf{P}(y|{\bm{W}}_{\star}^{T}\bm{\mathbf{x}})].$ (4)
Note that, in the limit $d\to\infty$, for $\bm{\mathbf{x}}\sim\mathsf{N}({\bm{0}},d^{-1}{\bm{I}})$, the Central Limit Theorem implies ${\bm{W}}_{\star}^{T}\bm{\mathbf{x}}\sim\mathsf{N}({\bm{0}},{\bm{I}}_{p})$ and $\mathsf{Z}(y)=\mathbb{E}_{\bm{\mathbf{z}}\sim\mathsf{N}({\bm{0}},{\bm{I}}_{p})}[\mathsf{P}(y|\bm{\mathbf{z}})]$.
@@
(Weak subspace recovery) Given an estimator $\bm{\mathbf{\hat{W}}}$ of $\bm{W}_{\star}$ with $\|\bm{\mathbf{\hat{W}}}\|_{\mathrm{F}}=\mathrm{\Theta}(d)$, we say we have weak recovery of a subspace $V\subset\mathbb{R}^{d}$, if
@@
Computational bottlenecks for weak recovery in the Gaussian multi-index models have been studied by Troiani et al. (2024) using an optimal generalized approximate message passing (GAMP) scheme, see Appendix A for a detailed discussion of the algorithm. In particular, for the appropriate choice of denoiser functions, given in eq. (58), AMP is provably optimal among first-order methods Celentano et al. (2020); Montanari and Wu (2024). Their work provides a classification of the directions in $\mathbb{R}^{p}$ in terms of computational complexity of their weak learnability. In particular, if and only if
@@
Troiani et al. (2024), Stability of the uninformed fixed point and critical sample complexity] If ${\bm{M}}={\bm{0}}\in\mathbb{R}^{p\times p}$ is a fixed point of the state evolution associated to the optimal GAMP (58), then it is an unstable fixed point if and only if $\norm{\mathcal{F}({\bm{M}})}_{F}>0$ and $n>\alpha_{c}d$, where the critical sample complexity $\alpha_{c}$ is:
@@
${\mathcal{F}}({\bm{M}})\coloneqq\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\left(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}{\mathrm{y}}]-{\bm{I}}\right)\bm{M}\left(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}{\mathrm{y}}]-{\bm{I}}\right)\right],$ (9)
@@
Moreover, if $\|\mathcal{F}({\bm{M}})\|_{\mathrm{F}}=0$, then ${\bm{M}}={\bm{0}}$ is a stable fixed point for any $n=\Theta(d)$.
The aim of our work is to close this gap, providing an estimation procedure that achieves weak recovery at the same critical threshold $\alpha_{c}$ defined in eq. (8), but crucially does not require an informed initialization.
@@
In what follows we restrict the problem defined in (3) to the set of link functions functions satisfying eq. (7). These functions (said to have a generative exponent $2$ in the terminology of Damian et al. (2024)) covers a large class of the relevant problems, with the exception of the really hard functions such as sparse parities, which cannot be solved efficiently with a linear (in $d$ number of samples Troiani et al. (2024).
Throughout this manuscript, we adopt the following definitions for tensor operations and tensor eigenpairs.
Definition 2.3.
(Tensor Contraction) For a tensor of order 4 of the form ${\underline{\bm{A}}}=(A_{ijkh})\in\mathbb{R}^{\nu\times p\times\nu\times p}$ with $\nu=n$ or $\nu=d$, we define the contraction of such tensor with a matrix $\bm{B}\in\mathbb{C}^{\nu\times p}$ by the operation ${\underline{\bm{A}}}\cdot\bm{B}\in\mathbb{R}^{d\times p}$ such that
$({\underline{\bm{A}}}\cdot\bm{B})_{kh}$ $:=\sum_{i\in\llbracket\nu\rrbracket}\sum_{j\in\llbracket p\rrbracket}A_{ijkh}B_{ij}\quad k\in\llbracket\nu\rrbracket,h\in\llbracket p\rrbracket.$ (10)
The tensor ${\underline{\bm{A}}}$ is symmetric if $\langle{\underline{\bm{A}}}\cdot\bm{B},\bm{C}\rangle=\langle\bm{B},{\underline{\bm{A}}}\cdot\bm{C}\rangle$ (equivalently $A_{ijkh}=A_{jikh}$ and $A_{ijkh}=A_{ijhk}$, for all $i,j\in\llbracket\nu\rrbracket$, $k,h\in\llbracket p\rrbracket$).
Definition 2.4.
(Tensor Inversion) A tensor ${\underline{\bm{A}}}$ is invertible if and only if there exists a tensor ${\underline{\bm{A}}}^{-1}\in\mathbb{R}^{\nu\times p\times\nu\times p}$ such that,
${\underline{\bm{A}}}^{-1}\cdot({\underline{\bm{A}}}\cdot\bm{B})={\underline{\bm{A}}}\cdot({\underline{\bm{A}}}^{-1}\cdot\bm{B})={\underline{\bm{I}}}\bm{B}=\bm{B}$ (11)
holds for all $\bm{B}\in\mathbb{\mathbb{missing}}{R}^{\nu\times p}$ and where ${\underline{\bm{I}}}=(I_{ijkh})$ is the identity tensor, $I_{ijkh}:=\delta_{ij}\delta_{kh}$.
In this framework, ${\underline{\bm{A}}}$ acts as a linear operator in the vector space of $\nu\times p$ matrices, with inner product $\langle\bm{B},\bm{C}\rangle\coloneqq{\rm Tr}(\bm{B}^{T}\bm{C})$ leading to the following natural definition of eigenpairs.
Definition 2.5.
(Tensor Eigenpairs) We say that $(\lambda,\bm{V})\in\mathbb{C}\times\mathbb{R}^{\nu\times p}\setminus\{\bm{0}_{\nu\times p}\}$ is an eigenpair of a tensor ${\underline{\bm{A}}}\in\mathbb{R}^{\nu\times p\times\nu\times p}$ if the following equation holds
${\underline{\bm{A}}}\cdot\bm{V}=\lambda\bm{V}.$ (12)
For the spectral theorem, a symmetric tensor admit an orthonormal basis of eigenvectors in $\mathbb{R}^{\nu\times p}$, corresponding to real eigenvalues.
Remark 2.6.
(Flattened Tensor) One can equivalently consider a flatten version of tensor, that is the matrix $\bm{flat}({\underline{\bm{A}}})\in\mathbb{R}^{\nu p\times\nu p}$ such that
$(\bm{flat}({\underline{\bm{A}}}))_{uv}:=A_{ijkh}$
$u=(i-1)p+j,\qquad v=(k-1)p+h\,.$ (13)
Similarly a flatten version of a matrix is defined as $\bm{flat}(\bm{V)}\in\mathbb{R}^{\nu p}$ such that $(\bm{flat}(\bm{V)})_{u}=V_{ij}$ with $u=(i-1)p+j$. A classical eigenpair of $\bm{flat}({\underline{\bm{A}}})$ is then related to the former definition 2.5 of the tensor eigenpair by $(\lambda,\bm{flat}(\bm{V}))$. By abuse of notation, we will write $\|{\underline{\bm{T}}}\|_{\mathrm{op}}=\|\bm{flat}({\underline{\bm{T}}})\|_{\mathrm{op}}$.
@@
We can now introduce the Spectral Methods we aim to investigate. Given a matix $\bm{\mathbf{X}}\in\mathbb{R}^{n\times d}$ with rows $\bm{\mathbf{x}}_{i}\sim\mathsf{N}({\bm{0}},d^{-1}{\bm{I}}_{d})$ and a vector of labels $\bm{\mathbf{y}}\in\mathbb{R}^{n}$ with elements sampled as in (3), define the tensor ${\underline{\bm{G}}}\in\mathbb{R}^{n\times p\times n\times p}$, acting on $n\times p$ matrices as described in Definition 2.5, as
$G_{ikjh}=\delta_{ij}\mathbb{E}[{\mathrm{z}}_{k}{\mathrm{z}}_{h}\big{|}{\mathrm{y}}_{i}]-\delta_{ij}\delta_{kh},\quad i,j\in\llbracket n\rrbracket,\;k,h\in\llbracket p\rrbracket.$ (14)
where $\delta_{ij}$ is the Kronecker delta.
Given these definitions, define the following two spectral estimators $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{L}}}}$,$\bm{\mathbf{\hat{W}}}_{{\underline{\bm{T}}}}$ of the weight matrix ${\bm{W}}_{\star}$ as
@@
    $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{L}}}}:=\sqrt{dN}\frac{\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\bm{\Omega}_{1}}{\|\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\bm{\Omega}_{1}\|_{\mathrm{F}}},$ (15)
    where $\bm{\Omega}_{1}$ is the (matrix) eigenvector corresponding to the eigenvalue $\gamma_{1}^{{\underline{\bm{L}}}}$ with largest real part of the tensor ${\underline{\bm{L}}}\in\mathbb{R}^{n\times p\times d\times p}$ defined for any $\bm{\Omega}\in\mathbb{R}^{n\times p}$ by:
    ${\underline{\bm{L}}}\cdot\bm{\Omega}\coloneqq(\bm{\mathbf{X}}\bm{\mathbf{X}}^{T}-{\bm{I}}_{n}){\underline{\bm{G}}}\cdot\bm{\Omega}\,.$ (16)
    Note that the constant $N$ can be chosen to fix the normalization of $d^{-1}\|\bm{\mathbf{\hat{W}}}\|_{\mathrm{F}}^{2}=N$.
@@
    $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{T}}}}:=\sqrt{dN}\frac{\bm{W}_{1}}{\|\bm{W}_{1}\|_{\mathrm{F}}},$ (17)
    where $\bm{W}_{1}$ is the (matrix) eigenvector associated to the largest eigenvalue $\gamma_{1}^{{\underline{\bm{T}}}}$ of the symmetric tensor map ${\underline{\bm{T}}}\in\mathbb{R}^{d\times p\times d\times p}$ defined as
    ${\underline{\bm{T}}}\cdot\bm{W}\coloneqq\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\left(({\underline{\bm{G}}}+{\underline{\bm{I}}})^{-1}\cdot(\bm{\mathbf{X}}\bm{W})\right)$ (18)
Note that the constant $N$ is arbitrary and can be chosen to fix $d^{-1}\|\bm{\mathbf{\hat{W}}}_{{\underline{\bm{L}}}}\|_{\mathrm{F}}^{2}=d^{-1}\|\bm{\mathbf{\hat{W}}}_{{\underline{\bm{T}}}}\|_{\mathrm{F}}^{2}=N$.

### v2

@@ 1 Introduction
A popular model to study learning problems in statistics, computer science and machine learning is the multi-index model, where the target function depends on a low-dimensional subspace of the covariates. In this problem, one aims at identifying the $p$-dimensional linear subspace spanned by a family of orthonormal vectors ${\bm{w}}_{\star,1},\dots,{\bm{w}}_{\star,p}\in\mathbb{R}^{d}$ from $n$ independent observations $({\bm{x}}_{i},y_{i})$ from the model:
@@
A classical problem in statistics (Friedman and Stuetzle, 1981; Yuan, 2011; Babichev and Bach, 2018), the multi-index model has recently gained in popularity in the machine learning community as a generative model for supervised learning data where the labels only depend on an underlying low-dimensional latent subspace of the covariates (Aubin et al., 2018; Damian et al., 2022; Dandi et al., 2024a).
Of particular interest to this work are spectral methods, which play a fundamental role in machine learning by offering an efficient and computationally tractable approach to extracting meaningful structure from high-dimensional noisy data. A paradigmatic example is the Baik–Ben Arous–Péché (BBP) transition (Baik et al., 2005), where the leading eigenvalue of a matrix correlates with the hidden signal — a phenomenon that is ubiquitous in machine learning theory. Beyond their practical utility, spectral methods often serve as a starting point for more advanced approaches, including iterative and nonlinear techniques. This leads us to the central question of this paper:
@@
While the optimal spectral method for single-index models are well understood (Luo et al., 2019; Mondelli and Montanari, 2018; Maillard et al., 2022; Mondelli et al., 2022), and their optimality in terms of weak recovery threshold established (Barbier et al., 2019; Mondelli and Montanari, 2018) their counterparts for multi-index models remain largely unexplored. This gap is particularly important, as multi-index models serve as a natural testbed for studying the computational complexity of feature learning in modern machine learning, and have attracted much attention recently (Abbe et al., 2023; Damian et al., 2022; Dandi et al., 2024a; Arnaboldi et al., 2023; Collins-Woodfin et al., 2024; Berthier et al., 2024; Simsek et al., 2024).
@@
In this work, we step up to this challenge by constructing optimal spectral methods for multi-index models. We introduce and analyze spectral algorithms based on a linearized message-passing framework, specifically tailored to this setting. Our main contribution is to present two such constructions, establish the reconstruction threshold for these methods and to show that they achieve the provably optimal threshold for weak recovery among efficient algorithms (Troiani et al., 2025).
@@
Recently, multi-index models have become a proxy model for studying non-convex optimization (Veiga et al., 2022; Arnaboldi et al., 2023; Collins-Woodfin et al., 2024). Ben Arous et al. (2021) has shown that the sample complexity of one-pass SGD for single-index model is governed by the information exponent of the link function. This analysis was generalized in several directions, such as to other architectures (Berthier et al., 2024), larger batch sizes (Arnaboldi et al., 2024a) and to overparametrised models (Arnaboldi et al., 2024c). A similar notion, known as the leap exponent, was introduced for multi-index models, where it was shown that different index subspaces might be learned hierarchically according to their interaction (Abbe et al., 2022, 2023; Bietti et al., 2023; Mousavi-Hosseini et al., 2024). The picture was found to be different for batch algorithms exploiting correlations in the data (Dandi et al., 2024b; Arnaboldi et al., 2024b; Lee et al., 2024), achieving a sample complexity closer to optimal first order methods (Barbier et al., 2019; Damian et al., 2024; Troiani et al., 2025). Troiani et al. (2025) in particular, provided optimal asymptotic thresholds for weak recovery within the class of first order methods.
@@
Spectral methods are widely employed as a warm start strategy for initializing other algorithms, in particular for iterative schemes (such as gradient descent) for which random initialization is a fixed point. Relevant to this work is the class of approximate message passing (AMP) algorithms, which have garnered significant interest in the high-dimensional statistics and machine learning communities over the past decade (Donoho et al., 2009; Bayati and Montanari, 2011; Rangan, 2011; Fletcher and Rangan, 2018). AMP for multi-index models was discussed in (Aubin et al., 2018; Troiani et al., 2025). Spectral initialization for AMP in the context of single-index models has been studied by (Mondelli and Venkataramanan, 2021).
@@
The interplay between AMP and spectral methods has been extensively studied in the literature, see for example (Saade et al., 2014; Lesieur et al., 2017; Aubin et al., 2019; Mondelli and Montanari, 2018; Mondelli et al., 2022; Maillard et al., 2022; Venkataramanan et al., 2022). In particular the idea of using message passing algorithm to derive spectral method has been very successful, leading to the non-backtracking matrix (Krzakala et al., 2013) and more recently to the Kikuchi hierarchy (Wein et al., 2019; Hsieh et al., 2023), and to non-linear optimal spectral methods for matrix and tensor factorization (Lesieur et al., 2017; Perry et al., 2018; Guionnet et al., 2023; Pak et al., 2024).
During the finalization of this work, an independent paper analysing a family of spectral estimators indexed by a pre-processing function $\mathcal{T}$ for Gaussian multi-index models Kovačević et al. (2025) appeared. Their results leverage tools from random matrix theory to characterize the asymptotic spectral distribution of this family, as well as the correlation of the top eigenvectors with the indices. They prove that a tailored $\mathcal{T}$ asymptotically achieves the optimal weak recovery threshold of Troiani et al. (2025) in the particular case where $\mathbf{G}(y)$ is jointly diagonalizable for all $y$, providing an alternative proof of 3.10 and Theorem 3.11 for this particular case.
@@ 2 Setting and Notations
To enhance readability, we adopt the following consistent notations through the paper: scalar quantities are denoted using non-bold font (e.g., $a$ or $A$), vectors are represented in bold lowercase letters (e.g., $\bm{a}$), and matrices are written in bold uppercase letters (e.g., $\bm{A}$). We further differentiate random variables depending on the noise with non-italic font (e.g, ${\mathrm{a}},\bm{\mathbf{a}},\bm{\mathbf{A}}$). We denote by $\langle\bm{a},\bm{b}\rangle$ the standard Euclidean scalar product, $\|\bm{a}\|$ the $\ell^{2}$-norm of a vector $\bm{a}$, $\|\bm{A}\|_{\mathrm{op}}$ the operator norm of a matrix $\bm{A}$ and $\|\bm{A}\|_{\mathrm{F}}$ its Frobenius norm. Given $n\in\mathbb{N}$, we use the shorthand $\llbracket n\rrbracket=\{1,\dots,n\}$. We denote $\mathbb{S}_{+}^{p}$ the cone of positive semi-definite $p\times p$ matrices. $(x)_{+}:=\max(0,x)$.
@@
and the labels’ marginal distribution $\mathsf{Z}(y)=\mathbb{E}_{{\bm{W}}_{\star}}\mathbb{E}_{\bm{\mathbf{x}}\sim\mathsf{N}({\bm{0}},d^{-1}{\bm{I}}_{d})}[\mathsf{P}({\mathrm{y}}=y|{\bm{W}}_{\star}^{T}\bm{\mathbf{x}})].$ Note that, in the limit $d\to\infty$, for $\bm{\mathbf{x}}\sim\mathsf{N}({\bm{0}},d^{-1}{\bm{I}})$, the Central Limit Theorem implies ${\bm{W}}_{\star}^{T}\bm{\mathbf{x}}\sim\mathsf{N}({\bm{0}},{\bm{I}}_{p})$ and $\mathsf{Z}(y)=\mathbb{E}_{\bm{\mathbf{z}}\sim\mathsf{N}({\bm{0}},{\bm{I}}_{p})}[\mathsf{P}({\mathrm{y}}=y|\bm{\mathbf{z}})]$.
@@
(Weak subspace recovery) Given an estimator $\bm{\mathbf{\hat{W}}}$ of $\bm{W}_{\star}$ with $\|\bm{\mathbf{\hat{W}}}\|_{\mathrm{F}}^{2}=\mathrm{\Theta}(d)$, we say we have weak recovery of a subspace $V\subset\mathbb{R}^{d}$, if
@@
Computational bottlenecks for weak recovery in the Gaussian multi-index models have been studied by Troiani et al. (2025) using an optimal generalized approximate message passing (GAMP) scheme, see Appendix A for a detailed discussion of the algorithm. In particular, for the appropriate choice of denoiser functions, given in eq. (40), AMP is provably optimal among first-order methods (Celentano et al., 2020; Montanari and Wu, 2024). Their work provides a classification of the directions in $\mathbb{R}^{p}$ in terms of computational complexity of their weak learnability. In particular, if and only if
@@
(Troiani et al., 2025), Stability of the uninformed fixed point and critical sample complexity] If ${\bm{M}}={\bm{0}}\in\mathbb{R}^{p\times p}$ is a fixed point of the state evolution associated to the optimal GAMP (40), then it is an unstable fixed point if and only if $\norm{\mathcal{F}({\bm{M}})}_{F}>0$ and $n>\alpha_{c}d$, where the critical sample complexity $\alpha_{c}$ is:
@@
${\mathcal{F}}({\bm{M}})\coloneqq\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[{\bm{G}}({\mathrm{y}})\bm{M}{\bm{G}}({\mathrm{y}})\right],$ (8)
@@
and ${\bm{G}}({\mathrm{y}})\coloneqq\mathbb{E}_{\bm{\mathbf{z}}\sim\mathsf{N}({\bm{0}},{\bm{I}})}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]$. Moreover, if $\|\mathcal{F}({\bm{M}})\|_{\mathrm{F}}=0$, then ${\bm{M}}={\bm{0}}$ is a stable fixed point for any $n=\Theta(d)$.
@@
The aim of our work is to close this gap, providing an estimation procedure that achieves weak recovery at the same critical threshold $\alpha_{c}$ defined in eq. (7), but crucially does not require an informed initialization. In what follows we restrict the problem defined in (3) to the set of link functions functions satisfying eq. (6). These functions (said to have a generative exponent $2$ in the terminology of Damian et al. (2024)) covers a large class of the relevant problems, with the exception of the really hard functions such as sparse parities, which cannot be solved efficiently with a linear (in $d$ number of samples (Troiani et al., 2025).
Troughout this manuscript we adopt the notation $\bm{a}={\rm vec}(\bm{A}\in\mathbb{R}^{b\times p})$, $b=n,d$, for the vector in $\mathbb{R}^{bp}$ with components $a_{(i\mu)}=A_{i\mu}$, $i\in\llbracket b\rrbracket,\,\mu\in\llbracket p\rrbracket$, where the double index $(i\mu)\in\llbracket b\rrbracket\times\llbracket p\rrbracket$ is a shorthand for the scalar index $i+(\mu-1)b$. Similarly, we say that ${\rm mat}(\bm{a})=\bm{A}\iff{\rm vec}(\bm{A})=\bm{a}$.
@@
We can now introduce the Spectral Methods we aim to investigate. Given a matix $\bm{\mathbf{X}}\in\mathbb{R}^{n\times d}$ with rows $\bm{\mathbf{x}}_{i}\sim\mathsf{N}({\bm{0}},d^{-1}{\bm{I}}_{d})$ and a vector of labels $\bm{\mathbf{y}}\in\mathbb{R}^{n}$ with elements sampled as in (3), define $\hat{{\bm{G}}}\in\mathbb{R}^{np\times np}$ as
$\hat{G}_{(i\mu),(j\nu)}\coloneqq\delta_{ij}G_{\mu\nu}({\mathrm{y}}_{i})=\delta_{ij}\mathbb{E}_{\bm{\mathbf{z}}\sim\mathsf{N}({\bm{0}},{\bm{I}})}[{\mathrm{z}}_{\mu}{\mathrm{z}}_{\nu}-\delta_{\mu\nu}|{\mathrm{y}}_{i}],$ (9)
and the following two spectral estimators $\bm{\mathbf{\hat{W}}}_{\bm{L}}$,$\bm{\mathbf{\hat{W}}}_{\bm{T}}$ of the weight matrix ${\bm{W}}_{\star}$ as
@@
    $\bm{\mathbf{\hat{W}}}_{\bm{L}}:=\sqrt{dN}\frac{\bm{\mathbf{X}}^{T}{\rm mat}\left(\hat{{\bm{G}}}\bm{\omega}_{1}\right)}{\|\bm{\mathbf{X}}^{T}{\rm mat}\left(\hat{{\bm{G}}}\bm{\omega}_{1}\right)\|_{\mathrm{F}}},$ (10)
    where $\bm{\omega}_{1}\in\mathbb{R}^{np}$ is the eigenvector corresponding to the eigenvalue $\gamma_{1}^{\bm{L}}$ with largest real part of the matrix $\bm{L}\in\mathbb{R}^{np\times np}$ defined as:
    $L_{(i\mu),(j\nu)}\coloneqq\left(\left[\bm{\mathbf{X}}\bm{\mathbf{X}}^{T}\right]_{ij}-\delta_{ij}\right)G_{\mu\nu}({\mathrm{y}}_{j})\quad i,j\in\llbracket n\rrbracket,\,\mu,\nu\in\llbracket p\rrbracket\,.$ (11)
@@
    $\bm{\mathbf{\hat{W}}}_{\bm{T}}:=\sqrt{dN}\frac{{\rm mat}({{\bm{w}}}_{1})}{\norm{{{\bm{w}}}_{1}}},$ (12)
    where ${\bm{w}}_{1}\in\mathbb{R}^{dp}$ is the eigenvector associated to the largest eigenvalue $\gamma_{1}^{\bm{T}}$ of the symmetric matrix $\bm{T}\in\mathbb{R}^{dp\times dp}$ defined as
    $T_{(k\mu),(h\nu)}\coloneqq\sum_{i\in\llbracket n\rrbracket}{\mathrm{X}}_{ik}{\mathrm{X}}_{ih}\left[{\bm{G}}({\mathrm{y}}_{i})\left({\bm{G}}({\mathrm{y}}_{i})+{\bm{I}}\right)^{-1}\right]_{\mu\nu}\quad k,h\in\llbracket d\rrbracket,\,\mu,\nu\in\llbracket p\rrbracket\,.$ (13)
Note that the constant $N$ is arbitrary and can be chosen to fix $d^{-1}\|\bm{\mathbf{\hat{W}}}_{\bm{L}}\|_{\mathrm{F}}^{2}=d^{-1}\|\bm{\mathbf{\hat{W}}}_{\bm{T}}\|_{\mathrm{F}}^{2}=N$.